OpenStudy (anonymous):

if 3x = 2(mod 7). find x

ganeshie8 (ganeshie8):

try, x = 1, 2, 3, 4, 5, 6

ganeshie8 (ganeshie8):

put x = 1 3*1 = 2 (mod 7) FALSE

ganeshie8 (ganeshie8):

put x = 2 3*2 = 2 (mod 7) FALSE

ganeshie8 (ganeshie8):

put x = 3 3*3 = 2 (mod 7) TRUE.

ganeshie8 (ganeshie8):

so, x = 3

ganeshie8 (ganeshie8):

you can always try like this, when the mod value is less

OpenStudy (anonymous):

actually whats mean 2(mod7) ?

OpenStudy (loser66):

the remainder of some number when divided by 7

ganeshie8 (ganeshie8):

2(mod 7) means, when divided by 7, the remainder is 2

OpenStudy (loser66):

@ganeshie8 Is there someway in general?

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

x = 2 (mod 7) means 7 | x - 2

OpenStudy (loser66):

by trying like that, we get just 1 solution, in fact, there are infinite

ganeshie8 (ganeshie8):

nope, there is oly one solution mod 7, since gcd(3, 7) = 1

ganeshie8 (ganeshie8):

3x = 2 (mod 7) solution is :- x = 3 (mod 7)

OpenStudy (loser66):

just discuss: as you say x = 3 (mod7) means x -3 = 7n , so, if the right hand side is 21, so, x = 18 if the right hand side is 35, so, x = 33........ so???

ganeshie8 (ganeshie8):

thats right ! and all satisfy x = 3 (mod 7) so we say there is a unique solution in mod7 world :)

OpenStudy (loser66):

I never took this course before, just read the book. That's why I am not sure about what I am saying , hehehe. thanks for explanation

ganeshie8 (ganeshie8):

np :) number theory is awesome. since im able to digest it, its not a big deal for u.. :)

OpenStudy (loser66):

thanks

OpenStudy (anonymous):

thanks @ganeshie8 ,, but im still confuse about this,, and I want to ask another question about this,, maybe after that, i can understand,, 1) X=5(mod 7) 2)X=2(mod 5)

ganeshie8 (ganeshie8):

good :) whats the question ?

ganeshie8 (ganeshie8):

let me give u a question :- 4x = 3 (mod 11)

ganeshie8 (ganeshie8):

find x

OpenStudy (anonymous):

okay go on,, and i already mention the question just now,,

ganeshie8 (ganeshie8):

1) X=5(mod 7) 2)X=2(mod 5) these are not questions :) these are answers ok :)

ganeshie8 (ganeshie8):

question is this :- 4x = 3 (mod 11) find x

ganeshie8 (ganeshie8):

4x = 3 (mod 11) what this guy actually asking is this :- find x such that 4x -3 is divisible by 11

ganeshie8 (ganeshie8):

can u find SOME x such that, 4x-3 is divisible by 11 ha ?

OpenStudy (anonymous):

ermmmm,,, hehehe,, dont know

ganeshie8 (ganeshie8):

tttrrrrrryyy :)

OpenStudy (loser66):

x = 5 (mod7) x =2 (mod5) solve for x , right? and they are a system, not separate, right?

ganeshie8 (ganeshie8):

yes, if thats one problem, then they are system. we can solve them using Chinese Remainder Theorem

OpenStudy (loser66):

yup, got you @ganeshie8 and I got the answer is x = 13 ( mod 35) hihihi

ganeshie8 (ganeshie8):

oww ! but im getting 12 (mod 35) :- x = 5 (mod 7) => x - 5 = 7k x = 5+ 7k x = 2 (mod 5) 5+7k = 2(mod 5) k = 1 (mod 5) k = 1 + 5m so, x = 5 + 7(1 + 5m) = 5 + 7 + 35m = 12 + 35m

OpenStudy (loser66):

@nadia.a do you want me guide you to understand how to solve his problem? 4x = 3 ( mod 11)?

OpenStudy (anonymous):

yessss

OpenStudy (loser66):

@ganeshie8 I will show you mine after guiding the asker, we must meet at that problem because I apply chinese theorem.

OpenStudy (loser66):

@nadia.a in Z11 , the remainder of a division must < 11, right?

ganeshie8 (ganeshie8):

sure :) chinese and commonsense both should give same answer :)

OpenStudy (loser66):

so, the remainder should be one of 1, 2, 3, 4, 5, 6, 7, 8, 9, 10 ,

OpenStudy (loser66):

the number is 4*x and we find out the remainder in Z11 is 3, now test 4*1 =4 and 11|4 = 0 remain 4 , got this or not?

OpenStudy (anonymous):

yes i got it,,

OpenStudy (loser66):

4*2 =8 and 11|8 = 0 remain 8 4*3 = 12 and 11|12 = 1 remain1 4*4 =16 and 11|16 = 1 remain 5 4*5 = 20 and 11|20 = 1 remain 9 4*6 = 24 and 11|24 = 2 remain 2 4*7 =28 and 11|28 = 2 remain 6 4*8 = 32 and 11| 32 = 2 remain 10 4*9 = 36 and 11|36 = 3 remain 3 bingo, we got remain 3 hehehe. so 4*9 =3 (mod11) compare to 4*x = 3 (mod11) we have 9 = x got me?

ganeshie8 (ganeshie8):

Loser you're very fast learner xD

OpenStudy (loser66):

Thank you!!!

ganeshie8 (ganeshie8):

wish my brain works like u when i touch new subjects -_-

OpenStudy (loser66):

@ganeshie8 it's not new to me, just because I learned by reading the book, I didn't have the confirmation from the prof. When seeing it, I would like to practice under the observance of an expert like you

ganeshie8 (ganeshie8):

ahh thats kind of u, but i get what u mean :)

OpenStudy (anonymous):

owhhhhhhhhhhhhhhh,, i see,, hahaha,, thanks both of u :D

OpenStudy (loser66):

ok, back to our problem, To Chinese theorem x =5 (mod 7) x =2 (mod 5) we have a = 2, b =4, m = 5 and n =7 so, x = a[n*n]+ b [m*1](mod m*n) x = 2[7*7] + 4 [5*1] (mod 35) x = 2 *49 + 20 (mod 35) x = 118 (mod 35) x = 13 (mod 35) @ganeshie8

ganeshie8 (ganeshie8):

okay, before looking at where it went wrong wid chinese, lets verify if the answer is correct or not :)

ganeshie8 (ganeshie8):

hypothesis :- x =5 (mod 7) x =2 (mod 5) Answer :- x = 13 (mod 35)

OpenStudy (loser66):

oh, yea, that 's the easiest way to confirm, ha!!

ganeshie8 (ganeshie8):

we ask two questions now for verifying the answer :- 1) Does 13 | 7 leave a remainder of of 5 ? 2) Does 13 | 5 leave a remainder of 2 ?

OpenStudy (loser66):

yup, I am wrong, hihihih... please, point it out.

ganeshie8 (ganeshie8):

you want the chinese way ha ?

OpenStudy (loser66):

aha

ganeshie8 (ganeshie8):

gimme 5 mnts to pull up chinese formula :)

OpenStudy (loser66):

@ganeshie8 I know what's wrong with me, b = 5 not 4

ganeshie8 (ganeshie8):

ack.... just a typo lol

OpenStudy (loser66):

Thaaaaaaaaaaaaanks a ton

ganeshie8 (ganeshie8):

x =5 (mod 7) x =2 (mod 5) solution = \(a_1 N_1x_1 + .... a_r N_r x_r (\mod 7 \times 5)\)

OpenStudy (anonymous):

if we add one more such as X=3(mod 5)

ganeshie8 (ganeshie8):

you want to solve the system :- x =5 (mod 7) x =2 (mod 5) x=3(mod 5) ?

ganeshie8 (ganeshie8):

the simple answer is NO solution, cuz you cannot have below simultaneously :- x =2 (mod 5) x=3(mod 5)

OpenStudy (loser66):

a little change x = 3 (mod 11)

ganeshie8 (ganeshie8):

x =5 (mod 7) x =2 (mod 5) x=3(mod 11)

ganeshie8 (ganeshie8):

you wanto use chinese / commonsense ? :)

OpenStudy (anonymous):

why NO ????

ganeshie8 (ganeshie8):

when you ask :- x =2 (mod 5) x=3(mod 5) what you're asking is : gimme a number that leaves remainders 2 and 3 when divided by 5

ganeshie8 (ganeshie8):

there is no single number wid that property. so NO solution.

OpenStudy (anonymous):

owh,, okay2,, i just noticed it :DD

ganeshie8 (ganeshie8):

good :)

ganeshie8 (ganeshie8):

you're also familiar wid chinese nadia ?

ganeshie8 (ganeshie8):

x =5 (mod 7) x =2 (mod 5) x=3(mod 11) solution :\( 5 N_1x_1 + 2 N_2 x_2 + 3 N_3 x_3 (\mod 7 \times 5 \times 11) \)

ganeshie8 (ganeshie8):

\(N_1 = 5\times 11\) \(N_2 = 7\times 11\) \(N_3 = 7\times 5\) \(N_1x_1 \equiv 1 (\mod 7)\) \(N_2 x_2 \equiv 1 (\mod 5)\) \(N_3x_3 \equiv 1 (\mod 11)\)

ganeshie8 (ganeshie8):

plug and chug. dont wry if they dint cover chinese yet, it wil come soon...

OpenStudy (anonymous):

@ganeshie8 i want to ask,, when i review back ur question 4X=3(mod 11),, i understand now,, so if i got that kind of question in exam, i have to try one by one the integer to find the remainder ?

ganeshie8 (ganeshie8):

thats one way, there is a standard way for solving linear congruences like this :- 4x = 3 (mod 11) => 4x -11y = 3 this is just a diaphontine equation. familiar wid solving diaphontine equations ? :)

OpenStudy (anonymous):

erm,, not really,, how ? tomorrow i hv exam about this topic,, help me :)

OpenStudy (anonymous):

@ganeshie8